# On Task Performance and Model Calibration with Supervised and Self-Ensembled In-Context Learning

### v2

@@ Abstract
Following the standard supervised fine-tuning (SFT) paradigm, in-context learning (ICL) has become an efficient approach propelled by the recent advancements in large language models (LLMs), yielding promising performance across various tasks in few-shot data setups. However, both paradigms are prone to suffer from the critical problem of overconfidence (i.e., miscalibration), especially in such limited data setups. In this work, we deliver an in-depth analysis of the behavior across different choices of learning methods from the perspective of both performance and calibration, as well as their interplay. Through extensive controlled experiments, we find that simultaneous gains for both task performance and calibration are difficult to achieve, and the problem of miscalibration exists across all learning methods in low-resource scenarios. To address this challenging trade-off between performance and calibration, we then investigate the potential of self-ensembling techniques applied at different modeling stages (e.g., variations of in-context examples or variations in prompts or different ensembling strategies). We justify the feasibility of self-ensembling on SFT in addition to ICL, to make the predictions more calibrated and have comparable or even better performance. Our work sheds light on which learning paradigm to choose and how to enhance both task performance and calibration of LLMs.
On Task Performance and Model Calibration with Supervised and Self-Ensembled In-Context Learning
[…]
@@ 1 Introduction
Machine learning and NLP have undergone a significant transformation recently, largely propelled by large language models (LLMs) Radford et al. (2019); Brown et al. (2020); Chowdhery et al. (2022); OpenAI (2023). Among different learning paradigms, Supervised Fine-Tuning (SFT) and In-Context Learning (ICL) have emerged as predominant methodologies Raffel et al. (2020); Dong et al. (2022), demonstrating commendable efficacy across many tasks. SFT tunes the model’s parameter and effectively specializes a (general-purpose) model to specific tasks by learning the knowledge in the training data and optimizing the objective. ICL, for each input, leverages the few-shot examples (i.e., the so-called demonstrations) to generate predictions without tuning model parameters and treating the model as a ‘black box’.
@@
Considering the different input format between training with SFT and inference with ICL, Min et al. (2022) and Chen et al. (2022) introduce the in-context examples into training phrase, which we call supervised in-context learning (SICL). However, when the demonstrations as a strong inductive bias get combined with SFT, it has been shown to become more likely to fall into the problem of overconfidence Desai and Durrett (2020); Jiang et al. (2021), and the predicted confidence distribution of ICL may be miscalibrated due to the bias in in-context examples Fei et al. (2023). Through our extensive experiments, we observe that both paradigms, SFT and ICL, suffer from the problem of miscalibration in low-resource scenarios.
@@
The important challenges of overconfidence and miscalibration, particularly in scenarios marked by limited data availability, underscore the need for a nuanced understanding of these paradigms. However, most of the previous work Mosbach et al. (2023); Sun et al. (2023) only focus on comparing solely the performance of SFT and ICL on out-of-distribution (OOD) data, targeting general-purpose LLMs. Here, we instead focus on studying task-specialized language models, where the behavior of different paradigms’ in-task performance along with their calibration remains an open research question. In addition, considering the possible issue of overconfidence and miscalibration, we pose and study another crucial research question: is it possible to ensure both in-task performance and well-calibrated LM behavior at the same time? Satisfying both requirements is critical to the application of the model in real-life setups: an applied system should provide both accurate and calibrated (trustworthy) predictions to be responsible.
@@
To address the challenges above, in this paper we first investigate the performance and calibration of different model-tuning and ICL methods, along with their interplay, on 7 classification datasets in limited data setups. Our empirical investigations unveil a task-dependant phenomenon of in-task performance and calibration with regard to whether the corpus has been seen by the model before, which also gains increasing attention over the research community in data contamination Zhu et al. (2023); Deng et al. (2023). We further observe that most of the results show relatively high calibration errors, which are in contrast with responsible LLM applications.
@@
In response to the possible miscalibration in low-resource settings, we introduce and explore the application of self-ensembling, inspired by the effect of ensembling of multiple independent models in improving the reliability of the final predictions Ovadia et al. (2019). We incorporate diverse variations in in-context examples and prompts, tailor them to different learning methods, and find that self-ensembling fortifies the model’s calibration without compromising performance. The overall flow of the self-ensembling method is illustrated in Figure 1. This strategic approach not only contributes to reducing the calibration error of model predictions by 43% on average but also streamlines resource utilization and time expenditures by focusing on modifications to a single model.
@@
Contributions. 1) We deliver a comprehensive empirical analysis with different choices of learning methods across a variety of tasks in limited data scenarios. 2) We show the task-dependent relationship between in-task performance and calibration of LLMs and provide practical guidelines for the choice of learning paradigms (§6.1). 3) We investigate and justify the feasibility of the self-ensembling method in enhancing both the performance and calibration of LLMs (§6.2). We release the code at https://github.com/cambridgeltl/ensembled-sicl.
@@ 2 Related Work
Learning Paradigms. Fine-tuning (FT) pretrained LMs has been used as an effective method to adapt them to specific tasks and datasets Devlin et al. (2019); Raffel et al. (2020). With the generation of much larger and more powerful LMs Brown et al. (2020); Chung et al. (2022), parameter-efficient fine-tuning (PEFT) He et al. (2022); Zhou et al. (2023b) has been proposed, where the central idea is to tune only a fraction of the model parameters to reduce computation and memory costs.
@@
Without tuning the model, in-context learning (ICL) has shown great potential, achieving promising performance on various tasks with demonstration examples Brown et al. (2020). Motivated by the positive results of ICL by concatenating multiple in-context examples to the LLM input at inference, Min et al. (2022) and Chen et al. (2022) introduce labeled in-context examples to the supervised training process (SICL). This has been further improved and utilized with pretraining Gu et al. (2023); Shi et al. (2023) and other training strategies Ye et al. (2023b, a); Wei et al. (2023).
@@
Sun et al. (2023) study task performance and stability with different PEFT methods in addition to ICL, including prompt tuning, and instruction tuning (IT) Singhal et al. (2022). Duan et al. (2023) explore the relationship between ICL and IT, and interpret ICL as implicit IT. Mosbach et al. (2023) compare the generalization ability of ICL and FT on out-of-distribution (OOD) data. Zhou et al. (2023c) show the superior performance of hard prompt tuning over standard FT in low-data scenarios. However, previous work has not explored all the learning methods systematically within low-data scenarios and has not investigated them through the joint optics of in-task performance, confidence, and their trade-offs.
@@
Calibrating LLMs. A well-calibrated LM should be accurate in terms of performance while also producing reliable confidence estimates for their predictions. When concatenating different in-context examples with various labels, tokens, and example ordering, ICL would be influenced by the bias in the concatenated examples Zhao et al. (2021), and various calibration methods have been proposed to mitigate this issue Wang et al. (2023a); Zhou et al. (2023a). Moreover, the model itself also contains certain ‘implicit’ label biases due to its (pre)training corpus, which would have an effect on the confidence estimation as well Fei et al. (2023). Furthermore, FT may suffer from miscalibration for both in-distribution and OOD data due to over-parameterization when adapting to the specific data Kong et al. (2020), and the miscalibration effect, as we investigate in this paper, might be even more pronounced in limited data scenarios.
@@
Ensembling Model Predictions has been used to mitigate the problem of overconfidence and improve the reliability of the final model predictions Ovadia et al. (2019). A standard ensembling practice is to train the model with different hyper-parameters or different initialization Wenzel et al. (2020); Lakshminarayanan et al. (2017). Concerning LLMs, Sun et al. (2022) ensemble fine-tuned LLMs to quantify the uncertainty with disagreement among different ensembling components. Gleave and Irving (2022) and Wang et al. (2023b) ensemble partially tuned LLMs, considering the computation resources for training and the storage for saving different LLM checkpoints. Although the proposed method achieves better and more reliable results, it takes a considerable amount of time, computation, and storage resources to train and save multiple models, which often makes them inapplicable to LLMs. In contrast to having several tuned models with supervised learning, Yao et al. (2023) demonstrate the feasibility of using self-ensembling in ICL without tuning the model. In this paper, we explore self-ensembling in the novel context of diverse learning paradigms and low-data setups and demonstrate that it is possible to improve model calibration without compromising performance.
@@ 3 Background: Learning Paradigms
In this paper, we analyze and compare four different learning paradigms in low-resource scenarios: zero-shot learning (ZSL), [^1] In the context of ZSL and later ICL there is no actual ’learning’ taking the place, and the model simply reacts to the provided prompt, but we have chosen the term ZSL for consistency with previous literature. in-context learning (ICL), supervised fine-tuning (SFT) and supervised in-context learning (SICL). With classification tasks in focus, we briefly describe each paradigm in what follows.
@@
Supervised Fine-Tuning (SFT). As mentioned, ZSL and ICL are inference-only paradigms treating the LM as a black box. On the other hand, FT first trains the model on the training set following the input format $f_{p}(x)$ from ZSL. Note that here we use SFT to refer to instruction-style fine-tuning with a prompting template. Inference with the tuned model $\mathcal{P^{\prime}}$ is then conducted in the same way as with ZSL. Both during training and inference, we can use different prompting templates to create variations in the model input, which we further elaborate on in §4.1.
@@
4 Methodology
There are two points that create possible variations that can be used for ensembling: 1) variation in the selection of in-context examples (for ICL and SICL), and 2) variation in the chosen prompt (for all the paradigms). Previous work focuses on 1) selecting a better combination of in-context examples Su et al. (2023) for the model or 2) generating an optimal prompting template Zhou et al. (2023c). On the other side, how the variation of multiple demonstration combinations and prompting templates influences the model behavior is still unexplored. Furthermore, we can 3) ‘self-ensemble’ the model based on different ensembling strategies. We now introduce all these variants.
4.1 Variation of Ensembling Components
Variation of Prompting Templates (Var-Prompt). Different prompting templates have shown high variance in task performance Mishra et al. (2022). By changing the wording in the templates $f_{p}$, we can also create variations even with the same input to the model. For each input $x$, we randomly select a prompting template $f_{p}^{\prime}$ from a set of available prompting template candidates. In ICL and SICL, the same template is also applied to the in-context examples, formatting the final input as $[f_{p}^{\prime}(x_{1}),y_{1};...;f_{p}^{\prime}(x_{M}),y_{M},f_{p}^{\prime}(x)]$. This makes it applicable not only to ICL and SICL, but also to ZSL and SFT as well.
4.2 Self-Ensembling Strategy
Majority Vote. We select the predicted results that have the highest accumulated probability across $K$ variants as the ensembling predictions. The accumulated probability $\mathcal{P}_{acc}$ for the predicted label $l_{i}$ is defined as follows:
$\mathcal{P}_{acc}(\hat{y}=l_{i})=\sum_{k=1}^{K}\mathcal{P}_{k}(\hat{y_{k}}=l_{i})\mathbb{I}(\hat{y_{k}}=l_{i}).$ (1)
We pick the variants that have the same prediction as the ensembling prediction and average the probability distribution of the selected components,
$\mathcal{P}_{ens}(y|x)=\frac{1}{K^{\prime}}\sum_{k=1}^{K^{\prime}}\mathcal{P}_{k}(y|x),$ (2)
where $K^{\prime}$ is the number of selected variants.
Mean Probability. We average the predicted probability distribution of $K$ variants and use the prediction that has the largest probability in the averaged distribution as the ensemble result $\hat{y_{ens}}$:
$\mathcal{P}_{ens}(y|x)=\frac{1}{K}\sum_{k=1}^{K}\mathcal{P}_{k}(y|x),$ (3)
[…]
Max Probability. For each possible value in the output space, we find the maximum probability of the predicted values across $K$ variants and use this as the prediction’s probability.
$\mathcal{P}^{\prime}(\hat{y}=l_{i}|x)=\max(\mathcal{P}_{j}(\hat{y}=l_{i}|x),j\in[1,K])$
Because the probability is obtained from different components, the summation of these probabilities is not guaranteed to be 1. Therefore, we apply the normalization on the new probability distribution: $\mathcal{P}_{ens}(y|x)={\rm Norm}(\mathcal{P}^{\prime}(y|x))$. The ensemble prediction is determined as the $\hat{y}$ that has the highest probability after the ensembling step.
4.3 Estimating Calibration
Beyond task performance of all the possible variants, we estimate the calibration of the model’s predictions (as a proxy towards model confidence) by using Expected Calibration Error (ECE) Guo et al. (2017). It divides the $n$ predicted results based on their confidence into $M$ bins $B_{1}$ to $B_{M}$ and then computes a weighted average over the absolute difference between the accuracy ${\rm acc}(B_{m})$ and mean confidence ${\rm conf}(B_{m})$ of the predictions within each bin. We set $M$ to 10 in this work.
@@ 4.3 Estimating Calibration
ECE measures the difference between the model’s empirical accuracy and its confidence (predicted probability). The smaller the ECE, the more confident the model prediction would be. We also report the negative log-likelihood (NLL) Hastie et al. (2001) $-\sum_{i=1}^{n}{\rm log}(p_{i})$ and information entropy (IE) $-\sum_{i=1}^{n}p_{i}{\rm log}(p_{i})$ as supplementary metrics of model’s (lack of) confidence.
5 Experimental Setup
Datasets and Evaluation Metrics. We consider 7 classification datasets that cover a range of label numbers and scenarios: SST-2, SST-5 Socher et al. (2013), RTE Wang et al. (2019), ANLI Nie et al. (2020), Measuring Hate Speech corpus Sachdeva et al. (2022), Intent Detection from NLU++ Casanueva et al. (2022) and Manifestos Lehmann et al. (2023). In order to simulate low-data setups, we sub-sample smaller training data from the full data for each dataset. The details of the datasets along with their corresponding evaluation metrics are provided in Appendix A.1.
Implementation Details. Unless noted otherwise, we use Flan-T5large Chung et al. (2022) as the main model in the experiments. Detailed training environments and hyper-parameters are provided in Appendix A.2 and A.3. Further, we provide all the prompting templates in Appendix D.1 and D.2. Regarding the self-ensembling experiments, for variations of in-context examples, we generate 20 different combinations of IC examples per each original training example. For variations of prompting templates, we manually create 4 different templates for each task. For Var-Both, we experiment with the same 4 templates and generate $4\times 5$ variants with different combinations of IC examples.
6 Results and Discussions
6.1 Comparison between Learning Methods

### v3

# Large Language Models are Miscalibrated In-Context Learners

@@ Abstract
When adapting ICL with or without fine-tuning, we are curious about whether the instruction-tuned language model is able to achieve well-calibrated results without suffering from the problem of overconfidence (i.e., miscalibration) considering its strong instruction following ability, especially in such limited data setups. In this work, we deliver an in-depth analysis of the behavior across different choices of learning methods from the perspective of both performance and calibration. Through extensive controlled experiments, we observe that the miscalibration problem exists across all learning methods in low-resource setups. To achieve simultaneous gain for both in-task performance and calibration, we then study the potential of self-ensembling applied at different modeling stages (e.g., variations of in-context examples or variations in prompts or different ensembling strategies) to make the predictions more calibrated and have comparable or even better performance. We find that self-ensembling with max probability produces robust and calibrated predictions. Our work reveals the potential calibration problem of using ICL despite the improvements in task performance and sheds light on which learning paradigm to choose. We also provide practical guidelines for choosing learning paradigms depending on whether the data has been seen by the model before and a worthwhile solution via self-ensembling on how to enhance both task performance and calibration of LMs, which we hope could encourage further study.
Large Language Models are Miscalibrated In-Context Learners
Chengzu Li1, Han Zhou1, Goran Glavaš2, Anna Korhonen1, Ivan Vulić1 1Language Technology Lab, University of Cambridge 2Center for Artificial Intelligence and Data Science, University of Würzburg {cl917, hz416, iv250, alk23}@cam.ac.uk goran.glavas@uni-wuerzburg.de
@@ 1 Introduction
Machine learning and NLP have undergone a significant transformation recently, largely propelled by language models (LMs) (Radford et al., 2019; Brown et al., 2020; Chowdhery et al., 2022; OpenAI, 2023). Among different learning paradigms, Supervised Fine-Tuning (SFT) and In-Context Learning (ICL) have emerged as predominant methodologies (Raffel et al., 2020; Dong et al., 2022), demonstrating commendable efficacy across many tasks. SFT tunes the model’s parameter and effectively specializes a (general-purpose) model to specific tasks by learning the knowledge in the training data and optimizing the objective. ICL, for each input, leverages the few-shot examples (i.e., the so-called demonstrations) to generate predictions without tuning model parameters and treating the model as a ‘black box’. Considering the different input format between training with SFT and inference with ICL, Min et al. (2022) and Chen et al. (2022) introduce the in-context examples into training phrase, which we call supervised in-context learning (SICL). However, when the demonstrations, as a strong inductive bias, get combined with SFT, it has been shown that LMs become more likely to fall into the problem of overconfidence (Desai and Durrett, 2020; Jiang et al., 2021); the predicted confidence distribution of ICL may be miscalibrated due to the bias in in-context examples (Fei et al., 2023). Through our extensive experiments, we observe that both paradigms, SFT and ICL, suffer from the problem of miscalibration in low-resource scenarios.
@@
The important challenges of overconfidence and miscalibration, particularly in scenarios marked by limited data availability, underscore the need for a nuanced understanding of these paradigms. These challenges could be more severe in instruction-tuned models considering their strong instruction-following abilities. However, most of the previous work (Mosbach et al., 2023; Sun et al., 2023) only focuses on comparing solely the performance of SFT and ICL on out-of-distribution (OOD) data, targeting general-purpose LMs. Here, we instead focus on studying instruction-tuned task-specialized language models, where the behavior of different paradigms’ in-task performance along with their calibration remains an open research question. Therefore, in this work, in addition to the task performance of the models, we are interested in: RQ1) how would ICL impact the calibration of LMs? Furthermore, considering the possible issue of overconfidence and miscalibration, we pose and study another crucial research question: RQ2) is it possible to ensure both in-task performance and well-calibrated LM behavior at the same time? Satisfying both requirements is critical to the application of the model in real-life setups: an applied system should provide both accurate and calibrated predictions to be responsible.
@@
To address the above challenges, we first investigate the performance and calibration of different model-tuning and ICL methods, along with their interplay, on 7 classification datasets in limited data setups. The experiments without ICL demonstrate that the vanilla language model is not necessarily calibrated despite of the high in-task performance it achieves. We find empirically that ICL doesn’t help to improve the calibration consistently across 5 of the 7 classification datasets. Our empirical investigations with ICL unveil a phenomenon between in-task performance and calibration, depending on whether the task dataset has been seen by the model before, which also gains increasing attention over the research community in data contamination (Zhu et al., 2023; Deng et al., 2023). We further observe that most of the results show relatively high calibration errors, which stands in contrast with the requirements of responsible LMs.
@@
Based on the findings above, we propose practical guidelines for choosing different learning paradigms, which suggests supervised tuning methods are likely to achieve better performance on unseen datasets while on seen datasets ICL combined with other ‘tweaks’ such as model calibration can be a better choice. In response to the possible miscalibration, we explore the application of self-ensembling as a potential solution (shown in Figure 1), inspired by the effect of ensembling of multiple independent models in improving the reliability of the predictions (Ovadia et al., 2019). We incorporate diverse variations in in-context examples and prompts, tailor them to different learning methods, and find that self-ensembling fortifies the model’s calibration by 43% on average without compromising task performance.
@@
Contributions. 1) We deliver a comprehensive empirical analysis with different choices of learning methods across a variety of tasks in limited data scenarios, demonstrating that ICL doesn’t improve the calibration of the model consistently (§5.1). 2) We show the relationship between in-task performance and calibration of LMs depending on whether the data has been seen by the model and provide practical guidelines for the choice of learning paradigms (§5.2). 3) We investigate and justify the feasibility of the self-ensembling pipeline in enhancing both the performance and calibration of LMs (§6). We release the code at https://github.com/cambridgeltl/ensembled-sicl.
@@ 2 Related Work
Learning Paradigms. Fine-tuning (FT) pretrained LMs has been used as an effective method to adapt them to specific tasks and datasets (Devlin et al., 2019; Raffel et al., 2020). With the generation of much larger and more powerful LMs (Brown et al., 2020; Chung et al., 2022), parameter-efficient fine-tuning (PEFT) (He et al., 2022; Zhou et al., 2024b) has been proposed, where the central idea is to tune only a fraction of the model parameters to reduce computation and memory costs.
@@
Without tuning the model, in-context learning (ICL) has shown great potential, achieving promising performance on various tasks with demonstration examples (Brown et al., 2020). Motivated by the positive results of ICL by concatenating multiple in-context (IC) examples to the LM input at inference, Min et al. (2022) and Chen et al. (2022) introduce labeled in-context examples to the supervised training process (SICL). This has been further improved and utilized with pretraining (Gu et al., 2023; Shi et al., 2023) and other training strategies (Ye et al., 2023; Wei et al., 2023).
@@
Sun et al. (2023) study task performance and stability with different PEFT methods in addition to ICL, including prompt tuning, and instruction tuning (IT) (Singhal et al., 2022). Duan et al. (2023) explore the relationship between ICL and IT, and interpret ICL as implicit IT. Mosbach et al. (2023) compare the generalization ability of ICL and FT on out-of-distribution (OOD) data. Zhou et al. (2023) show the superior performance of hard prompt tuning over standard FT in low-data scenarios. However, previous work has not explored all the learning methods systematically within low-data scenarios and has not investigated them through the joint optics of in-task performance, confidence, and their trade-offs.
@@
Calibrating LMs. A well-calibrated LM should be accurate in terms of performance while also producing reliable confidence estimates for their predictions. When concatenating different in-context examples with various labels, tokens, and example ordering, ICL would be influenced by the bias in the concatenated examples (Zhao et al., 2021), and various calibration methods have been proposed to mitigate this issue (Wang et al., 2023a; Zhou et al., 2024a; Liu et al., 2024). Moreover, the model itself also contains certain ‘implicit’ label biases due to its (pre)training corpus, which would have an effect on the confidence estimation as well (Fei et al., 2023). Furthermore, FT may suffer from miscalibration for both in-distribution and OOD data due to over-parameterization when adapting to the specific data (Kong et al., 2020), and the miscalibration effect, as we investigate in this paper, might be even more pronounced in limited data scenarios.
@@
Ensembling Model Predictions has been used to mitigate the problem of overconfidence and improve the reliability of the final model predictions (Ovadia et al., 2019). A standard ensembling practice is to train the model with different hyper-parameters or different initialization (Wenzel et al., 2020; Lakshminarayanan et al., 2017). Concerning LMs, Sun et al. (2022) ensemble fine-tuned LMs to quantify the uncertainty with disagreement among different ensembling components. Gleave and Irving (2022) and Wang et al. (2023b) ensemble partially tuned LMs, considering the computation resources for training and the storage for saving different LM checkpoints. Although the proposed method achieves better and more reliable results, it takes a considerable amount of time, computation, and storage resources to train and save multiple models, which often makes them inapplicable to LMs. In contrast to having several tuned models with supervised learning, Yao et al. (2023) demonstrate the feasibility of using self-ensembling in ICL without tuning the model. In this work, we explore self-ensembling in the novel context of diverse learning paradigms and low-data setups and show that it is possible to improve model calibration without compromising performance.
@@ 3 Background: Learning Paradigms
In this paper, we analyze and compare four different learning paradigms relevant to ICL in low-resource scenarios: zero-shot learning (ZSL), [^1] In ZSL and later ICL there is no actual ’learning’ taking place, and the model simply reacts to the provided prompt, but we have chosen the term ZSL for consistency with previous literature. in-context learning (ICL), supervised fine-tuning (SFT) and supervised in-context learning (SICL). With classification tasks in focus, we briefly describe each paradigm in what follows.
@@
Supervised Fine-Tuning (SFT). As mentioned, ZSL and ICL are inference-only paradigms treating the LM as a black box. On the other hand, SFT first trains the model on the training set following the input format $f_{p}(x)$ from ZSL. Note that here we use SFT to refer to instruction-style fine-tuning with a prompting template, which we see as the context. Inference with the tuned model $\mathcal{P^{\prime}}$ is then conducted in the same way as with ZSL. During training and inference, we can use different prompting templates to create variations in the model input, which we further elaborate on in §6.1.1.
@@
4 Experimental Setup
Datasets and Evaluation Metrics. We consider 7 classification datasets that cover a range of label numbers and scenarios: SST-2, SST-5 (Socher et al., 2013), RTE (Wang et al., 2019), ANLI (Nie et al., 2020), Measuring Hate Speech corpus (Sachdeva et al., 2022), Intent Detection from NLU++ (Casanueva et al., 2022) and Manifestos (Lehmann et al., 2023). In order to simulate low-data setups, we sub-sample smaller training data from the full data for each dataset. The details of the datasets along with their corresponding evaluation metrics are provided in Appendix A.1.
Implementation Details. Unless noted otherwise, we use Flan-T5large (Chung et al., 2022) as the main model in the experiments. Detailed training environments and hyper-parameters are provided in Appendix A.2 and A.3. Further, we provide all the prompting templates in Appendix D.1 and D.2.
Estimating Calibration. Beyond task performance of all the possible variants, we estimate the calibration of the model’s predictions (as a proxy towards model confidence) by using Expected Calibration Error (ECE) (Guo et al., 2017). It divides the $n$ predicted results based on their confidence into $M$ bins $B_{1}$ to $B_{M}$ and then computes a weighted average over the absolute difference between the accuracy ${\rm acc}(B_{m})$ and mean confidence ${\rm conf}(B_{m})$ of the predictions within each bin. We set $M$ to 10 in this work.
@@ 4 Experimental Setup
ECE measures the difference between the model’s empirical accuracy and its confidence (predicted probability). The smaller the ECE, the more confident the model prediction would be. We also report the negative log-likelihood (NLL) (Hastie et al., 2001) $-\sum_{i=1}^{n}{\rm log}(p_{i})$ and information entropy (IE) $-\sum_{i=1}^{n}p_{i}{\rm log}(p_{i})$ as supplementary metrics of model’s (lack of) confidence (Zhang et al., 2025). We define well-calibrated results by an ECE threshold below 0.2 to facilitate clearer discussion.
5 RQ1: How does ICL Impact Model Calibration?
In this section, we are interested in how ICL impacts the calibration of the language model. To answer this question, we experiment with ZSL, ICL, SFT and SICL to find out 1) whether the model is calibrated and 2) the differences in calibration with different learning methods. We first present a comprehensive analysis of various learning methods in low-resource settings, detailing model performance and calibration errors in Table 1. Full experimental results are shown in Appendix B.